# Ground Every Sentence: Improving Retrieval-Augmented LLMs with Interleaved Reference-Claim Generation

### v1

@@ 4.2 Evaluation Metrics
Building on our previous task definition, we focus on evaluating the model-generated outputs in three key areas:
@@
We first evaluate the quality of the answers generated by the model. We concatenate all claim parts in order to form the answer to the question, and follow the ALCE evaluation method to calculate the correctness and fluency of the answers. For answer correctness, in the ASQA dataset, we use Exact Match Recall (EM Rec.) to measure the percentage of golden short answers contained in the generated answers. In the ELI5 dataset, we adopt Claim Recall (Claim Rec.) to measure the percentage of key claims included in the answers. To evaluate answer fluency, we use MAUVE Pillutla et al. (2021) to measures the similarity between output and gold answer.
@@
Similar to ALCE, we employ the AutoAIS Bohnet et al. (2022) to measure the citation quality. Our citation quality is also measured by two metrics: 1) Correct Attribution Score (CAS), which determines if the answers is entirely supported by cited sentences; 2) Citation Redundancy Score (CRS), which identifies any redundant citation sentences.
For CAS, We use the TRUE Honovich et al. (2022) model to compute the entailment relationship between each reference part and the corresponding claim part. The final CAS score is the proportion of sentences predicted as correctly attributed among all the sentences in the answer. For CRS, since our method allows the model to select multiple contextual reference sentences for the same claim, we need to determine if the reference contains redundant sentences. The final CRS score is the proportion of non-redundant sentences relative to all sentences in the references.
Verifiability and Credibility
We employ two metrics to measure the Verifiability: 1) Citation Length, where shorter citation text typically reduces the time needed for fact-checking; 2) Attribution Ratio (AR), which represents the proportion of sentences in the output that are attributed.
We use Consistency Ratio (CR) to measure the Credibility, which determines the text consistency between the reference parts and the reference passages through string matching. Higher CR usually means the model generates fewer hallucinations when outputting the reference parts.
Method Model ASQA ELI5 Average
@@
ALCE VANILLA 66.8 40.4 73.6 72.5 57.2 12.0 51.1 50.0 62.0 26.2 62.4 61.3
SUMM 70.0 43.3 68.8 61.8 40.2 12.5 51.5 48.2 55.1 27.9 60.2 55.0
SNIPPET 69.8 41.4 65.3 57.4 62.9 14.3 50.4 45.0 66.4 27.9 57.9 51.2
ORACLE 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 61.9 35.1 66.2 64.4
0-shot GPT-4o 72.9 52.8 74.8 51.6 37.3 19.9 63.5 30.7 55.1 36.4 69.2 41.2
3-shot Vicuna-7B 81.4 50.5 74.0 70.0 38.6 13.9 73.2 59.7 60.0 32.2 73.6 64.9
Vicuna-13B 85.0 48.5 76.1 68.2 35.5 15.6 74.3 57.4 60.3 32.1 75.2 62.8
Llama3-8B 90.1 50.7 77.7 62.1 61.3 17.9 78.3 45.3 75.7 34.3 78.0 53.7
ChatGPT 74.9 52.6 72.5 63.4 27.8 17.8 68.6 50.8 51.4 35.2 70.6 57.1
GPT-4o 91.3 56.6 77.4 58.0 29.7 21.1 70.2 36.8 60.5 38.9 73.8 47.4
@@
Vicuna-13B 75.9 46.3 57.7 51.1 64.1 19.0 52.5 37.6 70.0 32.7 55.1 44.4
Llama3-8B 88.6 50.4 65.4 57.4 73.7 19.4 59.3 43.8 81.2 34.9 62.4 50.6
ReClaim w/IG Vicuna-13B 88.7 49.1 84.4 74.7 66.8 16.9 77.4 54.6 77.8 33.0 80.9 64.7
Llama3-8B 86.4 50.5 87.7 77.1 66.8 16.8 84.3 61.8 76.6 33.7 86.0 69.5
Table 2: Results on ALCE benchmark Gao et al. (2023b) and ReClaim. Definitions for Fluency, Correct. and Citation are in Section 4.2.
4.3 Methods and Baseline
To evaluate the effectiveness of the proposed method, we tested several models using various generation approaches. For each experimental setup, we used the ChatGPT results from the ALCE method as our baseline. Additionally, for the two test datasets, we adopted the Oracle-5 paragraphs provided by the ALCE as the reference passages in our model input.
Prompting
We directly prompt the model to generate answer text with citations. For GPT-4o OpenAI (2023), we employ both zero-shot and three-shot prompting Brown et al. (2020), while for GPT-3.5-turbo and open-source models such as Llama3-8B-Instruct Touvron et al. (2023) and the vicuna Chiang et al. (2023) series, we use 3-shot prompting. This approach allows us to comprehensively assess the performance and effectiveness of various models under different prompting scenarios.
Fine-tuned Model
[…]
As outlined in Section 3.4, we fine-tune the same model, generating two separate models: ReferModel and ClaimModel. We completed our training on four 80GB A800 GPUs. For the ReferModel, we performed full fine-tuning with a learning rate of 2e-5 over 3 epochs. For the ClaimModel, we employed Lora tuning Hu et al. (2021) with a learning rate of 5e-5 over 5 epochs. During the inference phase, we alternated between these two models to interleavingly generate the reference and claim parts. To prevent the model from generating overly short or long answers, we set constraints on the number of reference-claim pairs. For the ASQA dataset, the model must generate between 2 to 5 pairs. For the ELI5 dataset, due to its complexity, the model must generate between 4 to 6 pairs.
To investigate the roles of the two fine-tuned models, ReferModel and ClaimModel, in the ReClaim w/IG method, we conducted ablation experiments using the Llama3-8B-Instruct model. In these experiments, we alternately used only the fine-tuned ReferModel or ClaimModel, while employing the base model with 3-shot prompting to generate the other model’s output.
For these two fine-tuning methods, we adopted the constrained decoding approach described in Section 3.5 to limit the model’s output in the reference parts.
Method ASQA ELI5 Average
Fluency Correct. Citation Fluency Correct. Citation Fluency Correct. Citation Length
MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
ALCE 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 61.9 35.1 66.2 64.4 -
Citation-only 83.5 53.8 65.1 54.6 47.4 17.4 75.1 52.4 65.5 35.6 70.1 53.5 78.9
Claim-only 43.6 57.7 87.8 72.0 52.7 17.8 81.4 52.2 48.2 37.8 84.6 62.1 121.7
ReClaim w/IG 86.4 50.5 87.7 77.1 66.8 16.8 84.3 61.8 76.6 33.7 86.0 69.5 86.8
Table 3: The impact of different components under the ReClaim w/IG method on model generation results. Citation-only indicates using the fine-tuned ReferModel to generate references and the base model to generate claims. Conversely, Claim-only indicates using the fine-tuned ClaimModel to generate claims and the base model to generate references.
Method ASQA ELI5
Length Consistency Attri. Length Consistency Attri.
Citation Claim CR AR Citation Claim CR AR
ALCE 536.3 85.5 100.0 91.3 660.0 98.09 100.0 96.9
@@ 4.3 Methods and Baseline
[…]
@@
Table 4: The generated text length, consistency of references, and proportion of attributed answer sentences in different methods. CD indicates the use of constrained decoding, while UD indicates the absence of constrained decoding.
@@ 5.1 How to Improve the Quality of Answers and Citations?
The overall performance is presented in Table 2:
Prompting Works
Experimental results show that direct prompt models can achieve satisfactory outcomes. Compared to the ALCE method using Oracle-5 reference passages, our approach often matches or improves answer flunecy, correctness and citation correctness (CAS).
In all our prompting experiments, GPT-4o achieved the highest average correctness score (38.9) under 3-shot prompting. Additionally, Llama3-8B-Instruct achieved satisfactory results. Although it did not surpass the GPT model in terms of average correctness score, it demonstrated the best performance in fluency and CAS. Compared to the ALCE method, our model maintained response correctness while improving average fluency and CAS by 22.3% and 17.8%, respectively.
Although our method performs worse in CRS, the finer granularity of our citations minimizes the impact of redundant content. In contrast to paragraph-level citations, where each redundant citation introduces an entire paragraph of irrelevant content, our redundant citations only introduce a single sentence of irrelevant content. This does not significantly increase the fact-checking cost. Moreover, in many cases, redundant but continuous text helps in locating the citation in the original reference passages and aids in its understanding.
[…]
@@
Experimental results indicate that the ReClaim w/IG method outperforms other methods in two citation quality metrics while maintaining high fluency and correctness scores. Specifically, when using the Llama3-8B-Instruct model, our method outperforms the ALCE method with ChatGPT by 23.7%, 30.0%, and 7.9% in fluency, CAS, and CRS metrics, respectively, and only shows a 4.0% decrease in correctness.
[…]
As shown in Table 3, the experimental results demonstrate that the ReClaim w/IG method achieves the best performance in metrics of fluency and citation quality when compared to fine-tuning only the RefModel or the ClaimModel methods. This indicates the effectiveness of our fine-tuned dual-model approach. While the Claim-only method, which fine-tunes only the ClaimModel, yields the highest average answer correctness score (37.8), it comes at the cost of increased response length and reduced fluency. Additionally, numerous experiments indicate that improved citation quality often leads to a slight decrease in answer correctness. This is expected, as better attribution tends to constrain the source of information for model responses to the selected citations.
@@ 5.2 Can ReClaim Enhance the Verifiability and Credibility of RAG-based Question Answering?
To evaluate the verifiability of model responses and the impact of decoding constraints on their credibility, we measured the average length of citations, the consistency of citation texts (CR), and the proportion of sentences with attribution (AR) under different methods when using constrained decoding (CD) and unconstrained decoding (UD). The experimental results are shown in Table 4.
@@
Experimental results indicate that ReClaim’s average citations length is only about 20% of the length produced by the ALCE method, significantly reducing the fact-checking time cost. Compared to the ALCE method, our method ensures specific attribution for each response sentence, further improving the verifiability of model’s answer.
Decoding Constraints Improve Credibility
[…]
@@ 6 Conclusion
We propose a attributed text generation method, ReClaim, which adds sentence-level fine-grained citations to model-generated answer in RAG systems. This approach alternates between generating citations and answer sentences through Prompting or by fine-tuning the model.
The results show that our method improves citation quality while maintaining answer quality compared to the baseline method. Additionally, our approach significantly reduces the length of citations, thus decreasing the time cost required for fact-checking and further enhancing the verifiability of the model’s responses. Moreover, by using constrained decoding during citation generation, we ensure that each citation is composed of exact sentences from the source passages.
Although this paper focuses on long-form question answering task, we hope that our method can be extended to other tasks such as short-form question answering and Multi-document Summarization to improve the verifiability and credibility of model-generated responses.
@@
In this paper, our training dataset was exclusively targeted at long-form question-answering task, which reduces the generalization ability of our fine-tuning methods. Additionally, the construction of our training dataset did not account for the influence of irrelevant passages in the context on the model’s ability to find citations. This may weaken the model’s capability to filter out irrelevant information.
While our approach allows the model to synthesize information from multiple reference sentences for attribution, we did not specifically enhance this capability in our training data construction process and the model’s inference process.
@@
All original datasets used for training and testing were sourced from open and publicly accessible resources, and they are all approved for use in research purposes, thereby minimizing the risk of sensitive information leakage. While we employed LLMs for automated processing during the construction of training dataset, data cleansing was performed to prevent the introduction of additional noise and bias. We solely utilize the constructed dataset for model training. Although paragraph retrieval is not the focus of our work, information retrieved from large corpora may introduce some noise and bias into the responses generated by LLMs. To address these issues, we will optimize the data construction process and explore methods for retrieving noise-free and unbiased information.

### v2

@@ 4.2 Evaluation Metrics
Building on our previous task definition, we focus on evaluating the LLM-generated outputs in three key areas. Below, we introduce three evaluation dimensions, with more detailed information and calculation methods provided in Appendix C.3.
@@
We concatenate all claim parts in order to form the answer to the question, and follow the ALCE evaluation method to calculate the correctness and fluency of the answers. For answer correctness, in the ASQA dataset, we use Exact Match Recall (EM Rec.) to measure the percentage of golden short answers contained in the generated answers. In the ELI5 and EXPERTQA dataset, we adopt Claim Recall (Claim Rec.) to measure the percentage of key claims included in the answers. To evaluate answer fluency, we use MAUVE Pillutla et al. (2021) to measure the similarity between output and gold answer.
@@
Similar to ALCE, we employ the AutoAIS Bohnet et al. (2022) to measure the citation quality. Our citation quality is also measured by two metrics: 1) Correct Attribution Score (CAS), which determines whether the answer is entirely supported by cited sentences and is the most critical metric in our evaluation; 2) Citation Redundancy Score (CRS), which identifies any redundant citation sentences.
We use the NLI Honovich et al. (2022) model to compute the entailment relationship between each reference part and the corresponding claim part, and the final CAS score is the proportion of correctly attributed sentences in the answer.
We also use the NLI model need to determine if the reference contains redundant sentences, and the final CRS score is the proportion of non-redundant sentences relative to all sentences in the references.
Verifiability
We employ three metrics to measure the verifiability: 1) Citation length, where shorter citation text typically reduces the time needed for fact-checking; 2) Attribution Ratio (AR), which represents the proportion of sentences in the output that are attributed; 3) Consistency Ratio (CR), which represents the proportion of text consistency between the reference parts and the reference passages through string matching.
Method Model ASQA ELI5 EXPERTQA
@@
Prompting-Based
ALCE ChatGPT 64.4 48.9 74.5 72.7 59.4 21.3 57.8 56.0 53.7 20.5 66.7 64.9
Llama3-8B 79.2 55.2 54.7 54.6 45.4 20.5 42.8 39.3 56.7 20.2 51.4 47.8
w/Rerank 81.3 55.3 79.1 76.4 37.6 20.0 59.5 53.8 45.3 19.8 68.7 60.2
Post-hoc
ClosedBook Llama3-8B 50.5 28.9 13.7 13.7 66.8 16.8 17.4 17.4 37.6 25.4 11.2 11.2
Open-book Llama3-8B 73.7 53.2 52.1 52.1 26.6 20.8 31.9 32.0 58.7 21.0 36.9 36.8
Training-Based
Self-RAG Llama2-7B 70.6 38.7 53.3 66.2 33.1 9.7 23.1 33.9 19.0 12.5 9.4 12.1
RS+RL Llama2-7B 84.4 47.7 75.5 69.4 43.6 19.1 59.1 51.6 46.6 15.3 67.8 60.1
FRONT Llama2-7B 72.5 56.5 72.2 66.0 49.7 18.1 64.0 59.1 56.5 14.7 73.6 68.9
Our Methods
0-shot GPT-4o 72.9 52.8 74.8 51.6 37.3 19.9 63.5 30.7 59.4 18.5 73.4 26.5
@@
ChatGPT 74.9 52.6 72.5 63.4 27.8 17.8 68.6 50.8 30.6 14.1 72.7 55.7
GPT-4o 91.3 56.6 77.4 58.0 29.7 21.1 70.2 36.8 38.7 18.2 68.0 45.2
@@
Llama3-8B 88.1 53.5 92.1 86.1 71.6 17.8 89.9 67.5 63.5 14.0 90.1 68.6
Table 2: Results on ASQA, ELI5, EXPERTQA. Definitions for Fluency, Correct. and Citation are in Section 4.2.
4.3 Baselines
We use only methods that restrict citation sources to the top-k reference passages as our comparison baselines. We exclude methods that introduce additional retrieval steps, such as VTG Sun et al. (2023) and LLatrieval Li et al. (2024), from our comparison.
Prompting-Based
Following the ALCE Gao et al. (2023b) method, we prompt ChatGPT and Llama3.1-8B-Instruct with few-shot demonstrations that consist of a query, the top-5 retrieved passages, and an answer with inline citations. Additionally, we perform four samples of the generation from Llama3.1-8B-Instruct and select the best answer from each case as the experimental result for the w/Rerank method.
Post-hoc
Following the ALCE method, we generate answer using the Llama3-8B-Instruct model with the closed-book approach based on the given query. Then, for each statement in the answer, we search the top-100 passages for citations.
We also provide the LLMs with the top-5 reference passages, allowing it to generate the answer based on them and then re-identify citations within the top-5 passages. We refer to this method as open-book.
Training-Based
Following Self-RAG Asai et al. (2023), the LLM is trained to retrieve passages on demand, verify relevance, and generate answers based on the retrieved content. We directly provide the selfrag-7B (fine-tuned Llama2-7B model) with the top-5 passages for generation.
Following the method in Huang et al. (2024a), we use the LLM trained with fine-grained rewards (RS+RL, fine-tuned Llama2-7B model) to generate answer with citations. Additionally, we compare the fine-grained attribution method FRONT Huang et al. (2024b), which is fine-tuned based on the Llama2-7B model.
4.4 Methods
We evaluate ReClaim by testing several LLMs with different generation approaches.
ReClaim prompting
We directly prompt LLMs to generate answer with citations, using GPT models OpenAI (2023, 2022) and the Llama series Touvron et al. (2023) to evaluate the performance and effectiveness.
ReClaim ${}_{\texttt{Unified}}$
We follow the methodology in Section 3.3 and use the training dataset constructed in Section 3.2 to fully fine-tune the Llama3-8B-Instruct model, then conduct one-step generation with citations to evaluate the effectiveness of the ReClaim ${}_{\texttt{Unified}}$ method.
ReClaim w/IG
As outlined in Section 3.4, we fine-tune the same base model (Llama2-7B-hf and Llama3-8B-Instruct), generating two separate LLMs. During the inference phase, we alternate between these two LLMs to interleavingly generate the reference and claim parts.
For these two fine-tuning methods, we adopt the constrained decoding described in Section 3.5 to limit the generation of reference parts. For more experimental details, see Appendix C.
Method ASQA ELI5 EXPERTQA
Length Consistency Attri. Length Consistency Attri. Length Consistency Attri.
Citation Claim CR AR Citation Claim CR AR Citation Claim CR AR
ALCE 536.3 85.5 100.0 91.3 660.0 98.09 100.0 96.9 627.5 115.1 100.0 84.1
RS+RL 327.0 39.9 100.0 94.5 476.6 80.8 100.0 93.2 501.9 81.9 100.0 95.1
ReClaim ${}_{\texttt{3-shot}}$ 106.8 59.8 75.5 100.0 162.7 82.1 72.1 100.0 198.7 82.1 77.5 100.0
ReClaim ${}_{\texttt{Unified}}$ 77.9 52.9 100.0 100.0 139.8 93.1 100.0 100.0 150.9 99.2 100.0 100.0
ReClaim w/IG 82.8 68.9 100.0 100.0 145.0 104.7 100.0 100.0 157.5 109.1 100.0 100.0
@@ 4.4 Methods
Table 3: The generated text length, consistency of references, and proportion of attributed answer sentences in different methods. ReClaim ${}_{\texttt{3-shot}}$ does not use decoding constraints, while ReClaim ${}_{\texttt{Unified}}$ and ReClaim w/IG use decoding constraints.
Method ASQA ELI5 EXPERTQA
MAUVE EM Rec. CAS CRS MAUVE Claim Rec. CAS CRS MAUVE Claim Rec. CAS CRS
ReferModel-Only w/Extend 29.4 29.8 70.5 58.9 61.4 13.4 59.9 40.6 50.0 10.1 61.7 44.4
ReferModel-Only w/Sum 29.2 33.0 94.5 80.8 46.8 14.6 96.2 61.4 49.5 11.0 97.7 69.1
ClaimModel-Only 43.6 57.7 89.6 77.2 54.5 18.9 84.4 55.1 43.2 14.8 85.6 57.2
ReClaim w/IG 88.1 53.5 92.1 86.1 71.6 17.8 89.9 67.5 63.5 14.0 90.1 68.6
Table 4: Ablation study results. The underline indicates the second largest value.
4.5 Ablation Study
We primarily conduct ablation experiments on the ReClaim w/IG method to investigate the necessity of training data filtering and fine-tuning the ReferModel and ClaimModel. All fine-tuned models are based on the Llama3-8B-Instruct model.
Pre-Filtered Training Data
To investigate the necessity of using an NLI model for training data filtering, we fine-tune the LLM on the pre-filtered data while keeping the dataset size and training parameters unchanged. The experimental results are shown in Figure 3.
Figure 3: Comparison of the performance of LLMs trained on the training dataset before and after filtering on the ELI5 dataset. The comparative experimental results of the ASQA and EXPERTQA datasets are presented in Appendix Figure 9 and Figure 10.
The experimental results show that fine-tuning the LLM with data filtered by the NLI model improves its performance across various evaluation metrics.
Fine-tuned ReferModel Only
We use the fine-tuned ReferModel to generate reference parts and the base model to generate claim parts. When generating claims, it is essential to ensure that it adheres to the information in the preceding reference and maintains fluency with the previously generated claims. We explore two prompt strategies:
  1. 1.
    Extension: We provide the LLM with the previous reference and all preceding claims, requiring the LLM to extend the claims based on the information in the reference.
  2. 2.
    Summary: We provide the LLM with a reference so that it can directly generate a brief summary based on the previous reference.
Fine-tuned ClaimModel Only
We employ the 3-shot prompting approach to generate the reference section using the base model and utilize the fine-tuned ClaimModel for generating the claim section.
The experimental results are shown in Table 4.
4.6 Faithfulness Analysis
To further validate our approach in improving the credibility of model responses, we use GPT-4o-mini OpenAI (2023) as our evaluation model to evaluate the faithfulness metric of model answers against complete reference passages using the assessment method proposed by RAGAS Es et al. (2023). This helps us determine whether the LLM’s generated answers are fully based on the information in the given reference paragraphs.
We prioritize the CAS metric and use faithfulness as a supplementary measure to evaluate credibility improvements. Since faithfulness to a citation implies faithfulness to the original text, a high faithfulness score does not guarantee strong CAS performance, as an answer true to the full document may not align with the selected citation.
We evaluate the faithfulness metric on the test datasets, and the results for the ELI5 dataset are shown in Figure 4. Results for other datasets are shown in Appendix Figure 7 and 8.
Figure 4: The x-axis represents the accuracy of the LLM’s responses, while the y-axis shows the faithfulness score. For the Self-RAG and RS+RL methods, we use the fine-tuned 7B model, whereas for other methods, we employ Llama3-8B-Instruction as the base model.
Our method achieved the highest faithfulness score. This demonstrates that while our approach may slightly reduce answer quality, it significantly enhances answer faithfulness and minimizes unnecessary hallucinations linked to the LLM’s internal parameters.
The results show an inverse relationship between the accuracy of LLM-generated answers and their faithfulness to reference passages. Higher faithfulness and more granular citations have narrowed the scope of our answer sources, which may contribute to the lower accuracy in LFQA task.
@@ 5.1 How to Improve the Quality of Answers and Citations?
The overall performance is presented in Table 2.
ReClaim prompting Works
Experimental results show that directly prompting LLMs yields satisfactory outcomes. Our approach improves average answer fluency and citation accuracy (CAS) compared to baseline methods. Notably, Llama3-8B-Instruct surpasses other baselines in CAS scores, including ALCE+ChatGPT.
Although our method performs worse in CRS, the finer granularity of our citations reduces the impact of redundant content. Redundant citations only add a single irrelevant sentence, which does not significantly increase the cost of fact-checking.
ReClaim ${}_{\texttt{Unified}}$ Cannot Improve ATG
Experimental results show that the ReClaim ${}_{\texttt{Unified}}$ method significantly reduces citation quality (CAS). This indicates that fine-tuning the LLM in this way fails to teach it how to generate claims based solely on the information from the previous reference.
@@
Experimental results indicate that the ReClaim w/IG method outperforms other methods in two citation quality metrics while maintaining high fluency and correctness scores.
Compared to ALCE using ChatGPT, our method (using Llama3-8B) shows an average improvement of 31.3% on the CAS, 16.7% on the CRS, and 25.7% on the MAUVE across three test datasets, with only a 6.0% decrease in answer accuracy.
Specifically, we achieved an average CAS score of 90.7 across three test datasets, which is a crucial metric for assessing the degree of text attribution.
Compared to the ReClaim ${}_{\texttt{Unified}}$ method, the ReClaim w/IG approach’s biggest difference lies in the training and inference strategies during the claim generation phase. It filters out extraneous contextual information and trains the LLM to generate claims based solely on the preceding reference part. The significant improvements in citation quality demonstrate the effectiveness of the strategy adopted by the ReClaim w/IG method.
As shown in Table 4, the results of the ablation experiments indicate that fine-tuning two LLMs for alternating generation of references and claims achieves the most balanced performance.
While ReferModel-Only w/Sum method yields a high citation quality score, it compromises the accuracy and fluency of the answers.
On the other hand, using ClaimModel-Only method for generation achieves higher accuracy scores, but it negatively affects answer fluency, and the generation often tends to produce excessively long answers. Additionally, to ensure that the base model adheres to our format when generating reference parts, we need to provide multiple examples, which reduces the effective context window length.
@@ 5.2 Can ReClaim Enhance the Verifiability and Credibility of RAG-based Question Answering?
The results of the average length of citations, citation consistency (CR), and attribution ratio (AR) are shown in Table 3.
@@
Experimental results show that ReClaim’s average citation length is about 22% of the ALCE method, significantly reducing fact-checking time.
ReClaim ensures that each response sentence is supported by a specific citation source. By employing constrained decoding, our method ensures that every citation is directly extracted from the original text, significantly reducing the occurrence of hallucinations during the citation generation process. Together, these measures collectively enhance the verifiability of the generated answers.
ReClaim Improve Credibility
As mentioned in Section 5.1 regarding the improvement of attribution, this enhancement boosts the credibility of the LLM’s responses by providing a definitive source of evidence for the generated answers.
In addition, as shown in Figure 4, our method increases the faithfulness of the LLM’s responses to the overall reference passages by confining the sources of response information within the chosen citations. This improvement reduces the influence of the LLM’s internal knowledge on answer generation, thereby enhancing the credibility of our approach.
Our method strictly generates answers based on reference passages, aiming to improve citation quality in LLM responses through stringent constraints. The accuracy of our answers depends heavily on the information density of the reference passages and cannot rely on the LLM’s internal knowledge, which may also explain the slight decrease in answer accuracy in our approach.
@@ 6 Conclusion
We propose a attributed text generation method, ReClaim, which adds sentence-level fine-grained citations to LLM-generated answer in RAG systems. This approach alternates between generating citations and answer sentences, either through prompting or by fine-tuning LLMs.
The results show that our method improves citation quality while maintaining answer quality compared to the baseline methods. Additionally, our approach significantly reduces the length of citations, thus decreasing the time cost required for fact-checking and further enhancing the verifiability of the LLM’s responses. Moreover, by using constrained decoding during citation generation, we ensure that each citation is composed of exact sentences from the source passages.
@@
In this paper, our training dataset was exclusively targeted at long-form question-answering task, which reduces the generalization ability of our fine-tuning methods.
Additionally, our method requires explicitly outputting the cited sentences, which often leads to generating answers that are more than double the length. This results in increased output time for the LLM.
On one hand, while our approach allows the LLM to synthesize information from multiple reference sentences for attribution, we did not specifically enhance this capability during the training data construction and LLM inference processes. On the other hand, our answer template has certain limitations, as not every response generated by the LLM requires a citation from the reference passages. Therefore, our method has limitations in scenarios that require synthesizing information from multiple sources or necessitate multi-hop reasoning to draw conclusions.
@@
All original datasets used for training and testing were sourced from open and publicly accessible resources, and they are all approved for use in research purposes, thereby minimizing the risk of sensitive information leakage. While we employed LLMs for automated processing during the construction of training dataset, data cleansing was performed to prevent the introduction of additional noise and bias. We solely utilize the constructed dataset for model training. Although paragraph retrieval is not the focus of our work, the retrieved information from large corpora may introduce noise and bias into LLM-generated responses. To address these issues, we will optimize the data construction process and explore methods for retrieving noise-free and unbiased information.